OpenStudy (anonymous):

SHOW ALL OF THE STEP THAT YOU USE TO SOLVE THE PROBLEM 18.<1 and<2 are supplementary angles. m<1 is 4y+7,and m<2is 9y+4. FIND m<2

OpenStudy (mertsj):

What does supplementary mean?

OpenStudy (anonymous):

supplementary angles are two angles that equal 180 degrees

OpenStudy (mertsj):

So if angle 1 and angle 2 are supplementary, what does that mean?

OpenStudy (mertsj):

More precisely, supplementary angles are 2 angles whose SUM is 180

OpenStudy (mertsj):

how do you find a sum?

OpenStudy (mertsj):

What is the sum of 3 and 5 ?

OpenStudy (anonymous):

you add them so it be 8

OpenStudy (mertsj):

yes. you find a sum by adding. What is the sum of 4y and 9y?

OpenStudy (anonymous):

13 y

OpenStudy (mertsj):

What is the sum of 7 and 4?

OpenStudy (anonymous):

11

OpenStudy (mertsj):

What is the sum of 4y+7 and 9y+4

OpenStudy (anonymous):

13y+11

OpenStudy (mertsj):

So those two angles are supplementary. When you add them you get 180. So write the equation that says that.

OpenStudy (mertsj):

Are 120 and 60 supplementary angles?

OpenStudy (mertsj):

What is 120 + 60?

OpenStudy (anonymous):

180 so it right

OpenStudy (mertsj):

\[(4y+7)+(9y+4)=180\] \[13y+11=180\] \[13y=169\] \[y=13\] \[<2=9y+4=9(13)+4=121^{o}\]